What is the energy stored in the 1.00x10-5 F capacitor of a heart

defibrillator charged to 9.00x103 V?


1
Expert's answer
2022-05-25T10:27:26-0400
"W=\frac{CV^2}{2}"

"W=\frac{1.00*10^{-5}*(9.00*10^3)^2}{2}=405\:\rm J"